OpenStudy (anonymous):

A stone is dropped into a well in which the level of water is 'h' below the top of the well. If v is the velocity of sound. the time T after dropping the stone at which the splash is heard is ?

OpenStudy (anonymous):

It's easy though my question is would gravity act on sound , if no why not?

OpenStudy (anonymous):

@hartnn @Abhisar

OpenStudy (anonymous):

good question, but i assume it does not, since sound is a wave and travels through particles

OpenStudy (anonymous):

but it should act , as it would have some mass , and then we should apply kinematic equations for it too

OpenStudy (anonymous):

Options are a) T=2h/v b) T= Sq root 2h/g + h/v c) T= Sq.root 2h/g + h/2v d) T= Sq root hv/2g+ 2h/v

OpenStudy (anonymous):

So , i know (b) is correct , but still i had a query

OpenStudy (anonymous):

Let's take opinion of more people @myininaya

OpenStudy (anonymous):

sound in itself has no mass, it is carried by molecules that do have mass, but it's not affected like other object are by gravity. I'm not sure how well sound would travel in extreme gravity like neutron stars or the like, but under normal circumstances it doesn't really matter for sound if it travels away from earth or towards it, the speed will be the same.

OpenStudy (anonymous):

that was just a query though , btw even if it would act , it would negligible

OpenStudy (anonymous):

temperature is for more important than gravity, actually

OpenStudy (anonymous):

far*

OpenStudy (anonymous):

Light is actually found to bend under gravity.

OpenStudy (anonymous):

i think it's not the gravity bending the light, but rather spacetime itself is bent by gravity

OpenStudy (anonymous):

which makes it appear as if the light was bent, while it's in fact straight, but the space is not.

OpenStudy (anonymous):

oh maybe! :)